OpenStudy (anonymous):

MEDAL What is the oblique asymptote of the function f(x) = the quantity of x squared plus 6 x plus 10, all over x plus 2

OpenStudy (anonymous):

i think y=x-4

OpenStudy (anonymous):

we do you mean

OpenStudy (anonymous):

why is it important

OpenStudy (anonymous):

what about oblique

OpenStudy (anonymous):

but that is not one of my answer choices

OpenStudy (anonymous):

y=x+4 y=-x-4 y=x+1 y=x-1

OpenStudy (usukidoll):

what are the choices?

OpenStudy (usukidoll):

oh man I made an error....

OpenStudy (usukidoll):

you wouldn't mind if delete and start over do you?

OpenStudy (anonymous):

no

OpenStudy (usukidoll):

\[f(x) = \frac{x^2+6x+10}{x+2}\]

OpenStudy (usukidoll):